# Information Inequalities for Five Random Variables

### v1

@@ 13 $\langle 7,5,5\rangle$
We have also looked at how the newly discovered entropy inequalities delimit the $5$-variable entropy region. The triplet $\langle\alpha_{S},\beta_{S},\gamma_{S}\rangle$ yields the inequality
$(a,b\mkern 1.0mu{|}\mkern 1.0muz)+\alpha_{S}\big([abcd\mkern 1.5mu]+\mathcal{Z}\big)+\beta_{S}\mkern 1.5mu\mathcal{C}+\gamma_{S}\mkern 1.5mu\mathcal{D}\geqslant 0.$
Since the closure of the $5$-variable entropy region is a pointed convex cone, one can normalize it by assuming $(a,b\mkern 1.0mu{|}\mkern 1.0muz)=1$. An equivalent view is to take the cross-section of $\mkern 1.0mu\overline{\hbox{$\mkern-1.0mu\Gamma\mkern 4.0mu$}}\mkern-4.0mu^{*}_{\!5}$ by this hyperplane. Consider the $3$-dimensional subspace spanned by the vectors
@@
observe that $\mathbf{z}$ is negated. Normalize the $5$-variable entropic function $f$ so that it satisfies $f(a,b\mkern 1.0mu{|}\mkern 1.0muz)=1$, then project it to this subspace. Use the scalar products $\langle f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{x},f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{y},f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{z}\rangle$ as the projection coordinates. This $3$-dimensional cross-section of the $5$-variable entropy region is
$\Delta=\big\{\langle f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{x},f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{y},f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{z}\rangle\in\mathbb{R}^{3}:~~f\in\mkern 1.0mu\overline{\hbox{$\mkern-1.0mu\Gamma\mkern 4.0mu$}}\mkern-4.0mu^{*}_{\!5}\mbox{ such that }f(a,b\mkern 1.0mu{|}\mkern 1.0muz)=1\big\}.$
Clearly, points in $\Delta$ have non-negative $x$ and $y$ coordinates, while the $z$ coordinate can take both positive and negative values. Since $\mkern 1.0mu\overline{\hbox{$\mkern-1.0mu\Gamma\mkern 4.0mu$}}\mkern-4.0mu^{*}_{\!5}$ is a convex closed cone, $\Delta$ is both closed and convex. We concentrate on the part above the $xy$ plane:
@@
The Shannon inequalities provide no restriction whatsoever on $\Delta^{+}$ as any non-negative triplet $\langle x,y,z\rangle$ can be realized by some polymatroid. Points with non-negative $x$ and $y$ coordinates and $z=0$ are realized by linear polymatroids, thus the complete non-negative quadrant of the $xy$ plane is a part of $\Delta^{+}$. Our first non-Shannon inequality, generated by the triplet $S_{D_{1}}=\langle 1,0,0\rangle$, is
$(a,b\mkern 1.0mu{|}\mkern 1.0muz)+[abcd\mkern 1.5mu]+\mathcal{Z}\geqslant 0.$
@@
Refer to captionyyzzxx
Figure 4. Delimiting the $5$-variable entropy region
@@
Other extremal triplets provide additional linear constraints. Figure 4 illustrates the delimited part of the non-negative octant as viewed from the origin, and cut at $x\leqslant 2.5$ and at $y\leqslant 2.5$. The pictured bound of $\Delta^{+}$ is extended to larger values of $x$ and $y$. Along the $x$ and $y$ axes, this bound approaches the $xz$ and $yz$ coordinate planes as the functions $z=\sqrt{y}$ and $z=\sqrt{x}$, respectively. Along the $xy$ diagonal, the limiting behavior toward the $z$ axis is similar to the entropy function $z=-(x+y)\log(x+y)$. The corner point of the plateau $z=1$ has coordinates $\langle 1,1,1\rangle$. The $\theta(u\mkern 1.5mu{+}\mkern 1.5mu1,v\mkern 1.5mu{+}\mkern 1.5mu1)$ estimate gives a smooth bound on $\Delta^{+}$ which is asymptotically tight along the $x$ and $y$ axes.
@@ 9. Conclusions
Structural properties of the entropy region of four or more variables are mostly unknown. This region is bounded by linear inequalities corresponding to the non-negativity of Shannon information measures. Finding additional entropy inequalities is, and remains, an intriguing open problem. Previous works on generating and applying such non-Shannon entropy inequalities focused mainly on the four-variable case [4, 8, 15, 34], and only a few sporadic five-variable non-Shannon inequalities have been discovered [24]. This work provides infinitely many five-variable non-Shannon information inequalities by systematically exploring a special property of entropic vectors. Other works utilized the Copy Lemma, a method distilled from the original Zhang-Yeung construction by Dougherty et al. [15]. Our method is based on a different paradigm derived from the principle of maximum entropy, and is a special case of the Maximum Entropy Method described in[12]. As proven in Lemma 1, the principle of maximum entropy implies that every entropic polymatroid has an $n,m$-copy, which is a polymatroidal extension with special properties as defined in Definition 1. In Claim 2, we have proved that polymatroids having $n,m$-copies form a polyhedral cone, and hint at how its facets can be computed. Facet equations provide the potentially new non-Shannon entropy inequalities.
@@
Non-Shannon inequalities obtained from these computations are discussed in Section 6. Based on these experimental results, two infinite families of $5$-variable inequalities were defined. The first family in Theorem 11 is parametrized by downward closed subsets of non-negative lattice points. The second family in Theorem 16 has a single positive integer parameter. Inequalities in both families are proved to hold for polymatroids on $5$ elements that have an $n$-copy; consequently, they are all valid entropy inequalities. It is conjectured that they cover all inequalities that can be obtained by the applied method. In other words, if a polymatroid on $5$ elements satisfies all these inequalities, then it has an $n$-copy for all $n$. This conjecture is left as an open problem. The computational results confirmed this conjecture up to $n=9$.
Inequalities in the first family are investigated in Section 8 in more detail. These inequalities have the form
$(a,b\mkern 1.0mu{|}\mkern 1.0muz)+\alpha_{S}\big([abcd\mkern 1.5mu]+\mathcal{Z}\big)+\beta_{S}\mathcal{C}+\gamma_{S}\mathcal{D}\geqslant 0,$ (28)
@@
where the vector $\mathbf{v}_{S}=\langle\alpha_{S},\beta_{S},\gamma_{S}\rangle$ is computed from downward closed sets $S$ of nonnegative lattice points as specified in Definition 4. Such a vector $\mathbf{v}_{S}$ is extremal if the inequality (28) is not a consequence of other inequalities from the same family. Corollary 20 provides the following necessary condition for $\mathbf{v}_{S}$ to be extremal: $S$ should be either a horizontal or a vertical staircase. Based on additional properties of extremal vectors, proved in Claim 21, an incremental algorithm, sketched as Code 1 was used to generate all irreducible staircases, up to generation $60$. Extremal triplets are defined by irreducible staircases. The converse implication, which is valid for $n\leqslant 80$, is left as an open problem. Triplets ($\alpha_{S},\beta_{S},\gamma_{S}\rangle$ in the range $\beta_{S},\gamma_{S}\leqslant 20\alpha_{S}$, generated by irreducible staircases, are plotted in Figure 2. The number of new irreducible staircases that remained irreducible in the subsequent generation matches the sequence A103116 in the Encyclopedia of Integer Sequences [19]. It is an interesting open problem to prove the equality of these sequences.
@@
To illustrate how the newly discovered entropy inequalities delimit the $5$-variable entropy region, entropy vectors were normalized to satisfy $(a,b\mkern 1.0mu{|}\mkern 1.0muz)=1$ and projected onto a $3$-dimensional subspace. Part of the projection in the non-negative octant is denoted by $\Delta^{+}$. The Shannon inequalities do not provide any restriction on this part. Figure 4 illustrates the bounds implied by the new inequalities. While the non-negative quadrant of the $xy$ plane is known to be part of $\Delta^{+}$, and that it also contains points above that plane, it is an intriguing open problem whether our bound is, at least asymptotically, tight around the $x$ and $y$ axes. Showing that our bound is asymptotically tight at the point zero would amount to settling the long-standing open problem whether the entropic region is semi-algebraic.

### v2

@@ 8. The Minimal Set of Inequalities
We have also looked at how the newly discovered entropy inequalities delimit the five-variable entropy region. The triplet $\langle\alpha_{S},\beta_{S},\gamma_{S}\rangle$ yields the inequality
$(a,b\mkern 1.5mu{|}\mkern 1.5muz)+\alpha_{S}\big([abcd\mkern 1.5mu]+\mathcal{Z}\big)+\beta_{S}\mkern 1.5mu\mathcal{C}+\gamma_{S}\mkern 1.5mu\mathcal{D}\geqslant 0.$
Since the closure of the $5$-variable entropy region is a pointed convex cone, one can normalize it by assuming $(a,b\mkern 1.5mu{|}\mkern 1.5muz)=1$. An equivalent view is to take the cross-section of $\mkern 1.0mu\overline{\hbox{$\mkern-1.0mu\Gamma\mkern 4.0mu$}}\mkern-4.0mu^{*}_{\!5}$ by this hyperplane. Consider the three-dimensional subspace spanned by the vectors
@@
observe that $\mathbf{z}$ is negated. Normalize the five-variable entropic function $f$ so that it satisfies $f(a,b\mkern 1.5mu{|}\mkern 1.5muz)=1$, then project it to this subspace. Use the scalar products $\langle f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{x},f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{y},f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{z}\rangle$ as the projection coordinates. This three-dimensional cross-section of the five-variable entropy region is
$\Delta=\big\{\langle f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{x},f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{y},f\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{z}\rangle\in\mathbb{R}^{3}:~~f\in\mkern 1.0mu\overline{\hbox{$\mkern-1.0mu\Gamma\mkern 4.0mu$}}\mkern-4.0mu^{*}_{\!5}\mbox{ such that }f(a,b\mkern 1.5mu{|}\mkern 1.5muz)=1\big\}.$
Clearly, points in $\Delta$ have non-negative $x$ and $y$ coordinates, while the $z$ coordinate can take both positive and negative values. Since $\mkern 1.0mu\overline{\hbox{$\mkern-1.0mu\Gamma\mkern 4.0mu$}}\mkern-4.0mu^{*}_{\!5}$ is a closed convex cone, $\Delta$ is closed and convex. We concentrate on the part above the $xy$ plane:
@@
Shannon inequalities provide no restriction whatsoever on $\Delta^{+}$ as any non-negative coordinate triplet can be realized by some polymatroid. To show this, define $\mathbfit r_{I}$ for any subset $I$ of the ground set abcdzabcdz as
[…]
Note that $\lambda\mkern 1.5mu\mathbfit r_{I}$ is entropic for any non-negative $\lambda$. Let, moreover, $\mathbfit r^{\circ}$ be the function
[…]
as $A$ runs over the non-empty subsets of abcdzabcdz. Both $\mathbfit r_{I}$ and $\mathbfit r^{\circ}$ are exremal rays of $\Gamma_{\!5}$, so they are polymatroids. For arbitrary non-negative numbers $x,y,z$ the linear combination $f=\mathbfit r_{abcd}+x\mkern 1.5mu\mathbfit r_{ac}+y\mkern 1.5mu\mathbfit r_{ad}+z\mkern 1.5mu\mathbfit r^{\circ}$ satisfies $f(a,b\mkern 1.5mu{|}\mkern 1.5muz)=1$ and has coordinates $\langle x,y,z\rangle$, providing the required polymatroid.
Points of $\Delta^{+}$ with non-negative $x$ and $y$ coordinates and $z=0$ are realized by linear polymatroids; thus, the complete non-negative quadrant of the $xy$ plane is a part of $\Delta^{+}$. Our first non-Shannon inequality, generated by the triplet $S_{D_{1}}=\langle 1,0,0\rangle$, is
$(a,b\mkern 1.5mu{|}\mkern 1.5muz)+[abcd\mkern 1.5mu]+\mathcal{Z}\geqslant 0.$
@@
[…]
Figure 4. Points of $\Delta^{+}$ are on or below the indicated surface.
@@
Other extremal triplets provide additional linear constraints. Figure 4 illustrates the delimited part of the non-negative octant as viewed from the origin, cut at $x\leqslant 2.5$ and at $y\leqslant 2.5$. The depicted bound of $\Delta^{+}$ extends similarly to larger values of $x$ and $y$. Along the $x$ and $y$ axes, the bound approaches the $xz$ and $yz$ coordinate planes as the functions $z=\sqrt{y}$ and $z=\sqrt{x}$, respectively. Along the $xy$ diagonal, the limiting behavior toward the $z$ axis is similar to the entropy function $(x+y)\log(x+y)$. The corner point of the $z=1$ plateau has coordinates $\langle 1,1,1\rangle$. The $\theta(u\mkern 1.5mu{+}\mkern 1.5mu1,v\mkern 1.5mu{+}\mkern 1.5mu1)$ estimate provides a smooth version of this surface. Figure 5 plots the $0<x,y\leqslant 1$ part of this bound on a logarithmic scale. The flat “wings” around the $x$ and $y$ edges, rising at a slope of $1/2$, represent the region where the bound approaches the axes at the rate of $\sqrt{x}$ or $\sqrt{y}$, respectively. The diagonal hump of the surface rises as $t-\log t$, corresponding to the approximation $z\approx-(x+y)\log(x+y)$ as $x=y\to 0$.
Refer to caption $\scriptstyle{}-\log(x)$$\scriptstyle{}-\log(y)$$\scriptstyle{}-\log(z)$$\scriptstyle 0$$\scriptstyle 0$$\scriptstyle 7.5$$\scriptstyle 7.5$$\scriptstyle 15$$\scriptstyle 15$$\scriptstyle 22.5$$\scriptstyle 22.5$$\scriptstyle 5$$\scriptstyle 10$$\scriptstyle 15$$\scriptstyle 20$
Figure 5. Bounds of $\Delta^{+}$ around the origin on logarithmic scale
9. Matúš’ Inequalities
In [28], F. Matúš constructed an infinite family of five-variable non-Shannon inequalities. Similar to our bound of $\Delta^{+}$ in Figure 4, these inequalities provide the same quadratic outer bound on a special two-dimensional cross-section of the four-variable entropy region. By exhibiting examples for a quadratic inner bound in that cross-section, Matúš proved that the four-variable entropy region (and, consequently, the entropy region $\Gamma^{*}_{\!N}$ with $|N|\geqslant 4$) is not polyhedral. To describe his family of inequalities, let denote either $abcd\mkern 1.5mu]$ or $acbd\mkern 1.5mu]$.
Theorem 22 (F. Matúš [28]).
The following is an entropy inequality for every non-negative $k$:
[…]
Proof.
By induction on $k$. It clearly holds for $k=0$. To prove the inequality for $k\mkern 1.5mu{+}\mkern 1.5mu1$, suppose it holds for $k$ in every entropic polymatroid. Since the inequality to be proved does not contain entropy terms that intersect both $cd$ and $z$, the Maximum Entropy Principle implies that this inequality holds in every entropic polymatroid if and only if it holds in entropic polymatroids that additionally satisfy $(cd,z\mkern 1.5mu{|}\mkern 1.5muab)=0$; see also the proof of Lemma 17. Thus, without loss of generality, we may additionally assume that $cd$ and $z$ are conditionally independent given $ab$. By induction, (29) holds with $b$ and $c$ exchanged (as it holds in every entropic polymatroid); thus we have the following inequality:
[…]
The following inequalities hold in every polymatroid:
[…]
$(a,z\mkern 1.5mu{|}\mkern 1.5mub)+(a,b\mkern 1.5mu{|}\mkern 1.5muc)$ $\geqslant(a,z\mkern 1.5mu{|}\mkern 1.5muc)-(cd,z\mkern 1.5mu{|}\mkern 1.5muab),$
$(a,b\mkern 1.5mu{|}\mkern 1.5muz)+(a,c\mkern 1.5mu{|}\mkern 1.5mub)$ $\geqslant(a,c\mkern 1.5mu{|}\mkern 1.5muz)-(cd,z\mkern 1.5mu{|}\mkern 1.5muab).$
Adding the first line once, the second line, and the third line $k$ times to (30), and observing that $(cd,z\mkern 1.5mu{|}\mkern 1.5muab)=0$, yields the required inequality (29) for $k\mkern 1.5mu{+}\mkern 1.5mu1$. ∎
While Matúš’ inequalities do not appear in our families, they contain strikingly similar ones. Combining results from Theorem 11 and Theorem 16, for each non-negative $k$ the following is an entropy inequality:
[…]
As discussed at the end of Section 7, inequalities in (31) can also be proved by induction on $k$, similarly to the proof of Theorem 22. This means that they can be obtained by a different iterative method, exploiting the simple consequence of the Maximum Entropy Principle used above:
Suppose that the inequality $\psi$ on abcdzabcdz does not contain entropy terms that intersect both $cd$ and $z$. Then $\psi$ is a valid entropy inequality if and only if it holds for every entropic polymatroid that additionally satisfies $(cd,z\mkern 1.5mu{|}\mkern 1.5muab)=0$.
Since entropic polymatroids surely satisfy the Shannon inequalities, the first iteration extracts the additional inequalities that hold in polymatroids with $(cd,z\mkern 1.5mu{|}\mkern 1.5muab)=0$. This calculation has been done and provided two types of inequalities:
[…]
and their symmetric versions; see [10] or [26]. Consequently, they necessarily hold in entropic polymatroids. In the second iteration, inequalities from the first iteration are added to the Shannon inequalities, and then the consequences of this larger set are extracted, and so on.
Even completing the second iteration seems to be beyond reach since inequalities from the first iteration can be applied to any five subsets of the base set, resulting in an enormous set of inequalities. Restricting the substitution for permutations of the base elements, or certain one- and two-element sets, reduces the number of inequalities to a manageable level. For this and several other sparse sets, the computation has been done. The obtained new inequalities contained, as expected, the corresponding inequalities from (29) and (31); however, they also contained many others that were clearly not tight entropy inequalities: they were significantly weaker than those provided by Theorem 11. These preliminary results indicated no additional, structured family of non-Shannon inequalities provided by this alternate method.
@@ 10. Conclusions
Structural properties of the entropy region of four or more variables are mostly unknown. This region is bounded by linear inequalities corresponding to the non-negativity of Shannon information measures. Finding additional entropy inequalities is, and remains, an intriguing open problem. Previous works on generating and applying such non-Shannon entropy inequalities focused mainly on the four-variable case [4, 8, 15, 36], and only a few sporadic five-variable non-Shannon inequalities have been discovered [24]. This work provides infinitely many five-variable non-Shannon information inequalities by systematically exploring a special property of entropic vectors. Other works utilized the Copy Lemma, a method distilled from the original Zhang-Yeung construction by Dougherty et al. [15]. Our method is based on a different paradigm derived from the principle of maximum entropy and is a special case of the Maximum Entropy Method described in[12]. As proven in Lemma 1, the principle of maximum entropy implies that every entropic polymatroid has an $n,m$-copy, which is a polymatroidal extension with special properties as defined in Definition 1. In Claim 2, we have proved that polymatroids having $n,m$-copies form a polyhedral cone and hint at how its facets can be computed. Facet equations provide the potentially new non-Shannon entropy inequalities.
@@
Non-Shannon inequalities obtained from these computations are discussed in Section 6. Based on these experimental results, two infinite families of five-variable inequalities were defined. The first family in Theorem 11 is parametrized by downward closed subsets of non-negative lattice points. The second family in Theorem 16 has a single positive integer parameter. Inequalities in both families are proved to hold for polymatroids on five elements that have an $n$-copy; consequently, they are all valid entropy inequalities. It is conjectured that they cover all inequalities that can be obtained by the applied method. In other words, if a polymatroid on five elements satisfies all these inequalities, then it has an $n$-copy for all $n$. This conjecture is left as an open problem. The computational results confirmed this conjecture up to $n=9$.
@@
Inequalities in the first family are investigated in Section 8 in more detail. They are specified by triplets $\langle\alpha_{s},\beta_{s},\gamma_{s}\rangle$ determined by downward closed sets $s$ of nonnegative lattice points as discussed in Definition 4. Such a triplet is extremal if the corresponding inequality is not a consequence of other inequalities from the same family. Extremal triplets are determined by a special collection of downward closed sets called irreducible staircases. Based on the theoretical results in Corollary 20 and Claim 21, an incremental algorithm, sketched as Algorithm 1 was used to generate all irreducible staircases up to generation $60$. The converse implication, which is valid for the computed cases, that triplets generated by irreducible staircases are extremal, is left as an open problem. Triplets ($\alpha_{s},\beta_{s},\gamma_{s}\rangle$ in the range $\beta_{s},\gamma_{s}\leqslant 20\mkern 1.5mu\alpha_{s}$, generated by irreducible staircases, are plotted in Figure 2. The number of new irreducible staircases that remained irreducible in the subsequent generation matches the sequence A103116 in the Encyclopedia of Integer Sequences [31]. It is an interesting open problem to prove the equality of these sequences.
@@
To illustrate how the newly discovered entropy inequalities delimit the five-variable entropy region, entropy vectors were normalized to satisfy $(a,b\mkern 1.5mu{|}\mkern 1.5muz)=1$ and projected onto a three-dimensional subspace. Part of the projection in the non-negative octant is denoted by $\Delta^{+}$. The Shannon inequalities do not provide any restriction on this part. Figure 4 illustrates the bounds implied by the new inequalities. While the non-negative quadrant of the $xy$ plane is known to be part of $\Delta^{+}$, and that it also contains points above that plane, it is an intriguing open problem whether our bound is, at least asymptotically, tight around the $x$ and $y$ axes. Showing that our bound is asymptotically tight at the zero point would amount to settling the long-standing open problem of whether the entropic region is semi-algebraic.